TANGENTS AND NORMALS
at what points on the circle x^2 + y^2 – 2x – 4y + 1 = 0 the tangent is parallel to x axis 
 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Slope = 0
x^2 + y^2 - 2x - 4y - 1 = 0
2x + 2yy' - 2 - 4y' = 0
y’ = 0
2x - 2 = 0
x = 1
Put in curve:
1 + y^2 - 2 - 4y - 1 = 0
y^2 -4y - 2 = 0
y = \frac{4\pm \sqrt{24}}{2}
y = \frac{4\pm 2\sqrt{6}}{2}
y = 2\pm \sqrt{6}